# A Survey on Memory-Efficient Large-Scale Model Training in AI for Science

### v1

@@ II-B2 Medicine
LLMs have demonstrated significant potential in the medical and clinical fields. When trained on extensive medical datasets, these models can produce high-quality outputs, and support decision-making. The applications of such medical models are numerous, including medical knowledge retrieval, assistance in clinical diagnoses, consultations, and personalized medical services.
@@
Singhal et al. [3] found that the benchmark used to evaluate the clinical knowledge of the model is limited. To solve it, they proposed MultiMedQA, a benchmark composed of seven medical Q&A datasets. Further, they fine-tuned the command of PaLM [34] and got Flan-PaLM [32], which shows excellent performance in multiple-choice questions. Moreover, they suggested instruction prompt tuning, and the obtained Med-PaLM model can further adapt to medical field. The pre-training model of Med-PaLM, PaLM, uses ZeRO stage 3 and TP to shard model parameters across all TPUs during the training process and uses gradient checkpointing to reduce the memory overhead of activations.
@@
Tu et al. [5] aim to further explore the potential of generalist models (those capable of performing multiple tasks without fine-tuning, such as GPT-3, PaLM, and GPT-4) in medical field. Utilizing the language and multimodal foundation models PaLM 540B and PaLM-E 562B [55], the researchers develop Med-PaLM M. They also designed a multimodal biomedical benchmark called MultiMedBench. As a large-scale biomedical generalist model, Med-PaLM M demonstrates strong performance across all tasks on MultiMedBench, matching or exceeding the capabilities of task-specific models.
@@
Currently, most LLMs utilized in medical field are either closed source or have relatively small model sizes. Chen et al. [58] believe that while models designed for generalist tasks typically perform well across tasks, their performance in certain tasks falls short compared to models tailored for those specific purposes. They developed Meditron-70B through continued pretraining of Llama-2 and fine-tuned it for particular tasks, resulting in superior performance compared to all open-source generalist and medical LLMs across all medical benchmarks. Furthermore, to enable training, they extended Nvidia’s Megatron-LM [76] and developed Megatron-LLM [^2] https://github.com/epfLLM/megatron-LLM, a distributed training library that accommodates various model architectures. They also implemented 3D parallelism to optimize the training efficiency of models at this scale.
@@
HuatuoGPT [59] aims to assist doctors with diagnosis and treatment in Chinese scenarios, it combines the dialogue capabilities of ChatGPT with the expertise of real-world physicians. By employing instruction fine-tuning and reinforcement learning techniques, it exploits distilled instruction data generated by ChatGPT alongside actual response data from doctors. The upgraded version, HuatuoGPT-II [61], successfully passed the Chinese National Pharmacist Licensure Examination in October 2023 and nearly all medical qualification exams, highlighting its robust capabilities within the Chinese medical landscape. ZeRO data parallelism is enabled during the training of HuatuoGPT and HuatuoGPT-II.
II-B3 Biomedicine
@@
II-B4 Chemistry
The application of LLM in the field of chemistry covers multiple aspects, including molecular design, molecular property prediction, and chemical reaction prediction. Additionally, LLMs can be used to analyze and summarize chemical literature, providing researchers with efficient knowledge retrieval and decision-making support. Their powerful sequence processing capabilities enable the handling of complex chemical terminology and diverse data formats, thereby fostering advancements in automation and intelligence in the field of chemistry.
@@ II-B4 Chemistry
Irwin et al. pre-trained BART [68] on a dataset containing a large number of simplified molecular line entry system (SMILES) strings to capture the structural relationships and potential properties between chemical molecules, resulting in the Chemformer model [9]. Chemformer is capable of both sequence-to-sequence tasks (e.g. reaction prediction, molecular optimization), and discriminative tasks (e.g. molecular property prediction, biological activity analysis). It achieves or exceeds the performance of state-of-the-art models on multiple benchmarks while significantly reducing convergence time. This effectively addresses the challenges posed by limited computing resources and time constraints.
@@
The objective of molecular design and generation tasks is to create new molecules with specific functions or properties for applications such as drug development, materials science, and catalyst development. Traditional methodologies depend heavily on the expertise of specialists and costly computational simulations. However, the introduction of LLMs has opened new possibilities in this field. Fang et al. introduced a molecular generative pre-trained language model called MolGen [7], which has enhanced its comprehension of chemical structures and semantics through a two-stage pre-training strategy. This strategy encompasses molecular syntax and semantic learning, utilizing self-sacrificing embedded strings (SELFIES) along with domain-agnostic prefix tuning. Furthermore, the study proposed a chemical feedback paradigm that aligns generated molecules with real-world chemical preferences by adjusting their probability distribution, effectively reducing the issue of ”molecular hallucination” (i.e. generated molecules lacking expected functionality in practical applications). Frey et al. [6] conducted an in-depth study on the neural scaling behavior of deep chemical models in relation to dataset and model size, aiming to explore how model performance is enhanced with increased resource investment. They designed and trained a generative pre-trained Transformer model named ChemGPT with over 1B parameters for small molecule generation. By varying both the model and dataset sizes by several orders of magnitude, they demonstrated an improvement in pre-training loss performance as the scale increased.
II-B5 Meteorology
@@
Weather forecasting represents a crucial application scenario in meteorology, providing the ability to predict future weather changes, which benefits daily life, agriculture, transportation, and other fields. Over the past decade, the swift advancement of high-performance computing devices has led to significant progress in numerical weather prediction (NWP), enhancing the accuracy of forecasts for daily weather, extreme weather events, and climate change. However, as the growth of computing power slows and the complexity of physical models increases, the limitations of traditional methods have become more pronounced. Recently, the rapid progress in deep learning has emerged as a promising avenue for addressing these challenges in the field.
@@
The Pangu-Weather model [10], [11], proposed by Bi et al., employs the 3D Earth-Specific Transformer (3DEST) architecture to process 3D climate data, which is a variant of Swin Transformer [70], [71]. It is the first AI-driven approach to surpass the forecast accuracy of NWP methods. Notably, Pangu-Weather achieves an inference speed of just 1.4 seconds on a single GPU, making it 10000× faster than the previous most accurate NWP method (operational IFS [80]). While the accuracy of 3D models can surpass 2D models, this enhancement comes with a considerable increase in memory requirements, leading to limited network depth. The study employs DP to train on 192 GPUs, with a batch size of 1 assigned to each GPU.
@@
The inherent uncertainty in weather forecasting is inevitable, and the degree of this uncertainty, along with cumulative errors, increase as the forecast period lengthens. In response to this challenge, Chen et al. introduced FuXi [12], a cascaded model architecture based on U-Transformer, which effectively reduce cumulative errors and improves both forecast accuracy and duration. The model pre-training was conducted on a cluster of 8 Nvidia A100 GPUs, utilizing techniques such as FSDP [81], BF16 mixed precision training, and gradient checking to optimize memory usage.
@@
Inspired by the concept of foundational models [82], Nguyen et al. designed and trained a basic model named ClimaX [13], aimed at effectively adapting to general tasks related to the Earth’s atmosphere. ClimaX can be trained on heterogeneous datasets to solve diverse climate and weather tasks. Remarkably, even with pre-training conducted at lower resolutions and with limited computational resources, ClimaX surpasses previous data-driven models in weather and climate prediction benchmarks. The pre-training used 80 NVIDIA V100 32GB GPUs and enabled FP16 mixed precision training. Subsequently, Bodnar et al. introduced Aurora [14], a large-scale atmospheric foundation model designed to excel in various prediction tasks, particularly in regions with limited data or extreme weather conditions. After training on diverse weather and climate datasets, Aurora can effectively adapt to new atmospheric prediction tasks by learning a universal representation of atmospheric dynamics. The model was trained on 32 GPUs with 1 batch size per GPU. To meet memory constraints, BF16 mixed precision training and gradient checkpointing were also used.
@@
Despite the progress achieved in applying large-scale models across various domains, the continuous increase in memory overhead during model training remains a pressing concern that requires attention. We identify the underlying causes of this issue from the following key factors:
@@
  • The demand for handling diverse data structures. In scientific fields, the use of high-resolution images and multimodal data significantly contributes to increased memory consumption during large model training. High-resolution images contain a substantial amount of pixel information, thereby amplifying the demands of data processing and storage. Additionally, multimodal data requires the handling of diverse data types, such as images, text, and audio. Each modality necessitates an independent processing pathway and fusion mechanism, which further elevates the complexity and memory requirements of the model.
III Memory-Efficient Training Techniques of Transformers
LLMs are evolving rapidly and find extensive applications across various domains. However, the training of these large-scale models imposes significant demands on computing power and memory of computing devices. In this context, it becomes essential to explore memory-efficient strategies for training large models. These strategies focus on optimizing memory utilization and making more efficient use of computing resources. While LLMs have gained considerable popularity in scientific fields and have produced noteworthy results, many works have adopted no or few optimizations to improve training efficiency and reduce memory overhead. In the rest of this section, we will delve into various memory-efficient training techniques specific to transformer-based LLMs.
III-A Distributed Training
As the scale of models continues to increase, a single GPU’s memory can no longer accommodate complete model parameters, optimizer states, and intermediate activations. Traditional training strategies for large-scale models may take weeks or even months. However, distributed training allows the model to be divided across multiple devices, enabling the training of extremely large models with limited single GPU memory and significantly reducing training time.
III-A1 Data Parallelism
(a) Parameter Server
(b) Decentralized
Figure 2: Different architectures for data parallelism.
The memory overhead during model training primarily comes from model states and intermediate activations. The model states can be further divided into optimizer states, gradients, and parameters. Assuming the number of parameters in the model is denoted as $P$, and utilizing Adam optimizer with FP16 mixed precision during training, the memory overhead of model parameters is $2P+4P=6P$ bytes. The memory overhead of gradients is $2P$ bytes, and optimizer states consume $4P+4P=8P$ bytes. Thus, the total memory overhead is $16P$ bytes. If training GPT-3 (175B), would require approximately 2.8TB of memory. To address the redundancy of model states, Rajbhandari et al. introduced Zero Redundancy Optimizer (ZeRO) [75], which shards redundant model states across devices, reducing memory overhead up to $\frac{1}{N_{d}}$, where $N_{d}$ represents the number of devices. The core principle of ZeRO is that each device maintains only a fraction of the model states instead of a complete copy. The implementation of ZeRO has been integrated into Microsoft’s open-source distributed training framework, DeepSpeed [85], making it easily accessible for researchers and developers.
Inspired by ZeRO, PyTorch developed Fully Sharded Data Parallel (FSDP) [81]. A report from Hugging Face [^3] https://huggingface.co/blog/deepspeed-to-fsdp-and-back indicates that the performance of PyTorch FSDP is comparable to that of DeepSpeed ZeRO. Since PyTorch v2.4444https://github.com/pytorch/pytorch/releases/tag/v2.4.0, the PyTorch team has restructured FSDP and introduced FSDP2, adding additional optimizations to enhance performance.
Figure 3: Splitting of matrix multiplication.
III-A2 Tensor Parallelism
The rapid increase in model parameters has rendered the memory and computing resources of a single device insufficient to support the training of the entire model. Although ZeRO can partition the model across computing devices, it does not reduce computational workloads on each device. For large models with vast parameters and high computational demands, relying solely on DP becomes inadequate for meeting the training needs. Model parallelism (MP) has emerged as a crucial technique for further improving the scale of deep learning models. MP primarily includes two techniques: tensor parallelism (TP) and pipeline parallelism (PP), which will be discussed in this section and the following section. TP, also referred to as intra-layer parallelism, finely divides the computation within each model layer into smaller tensor operations that can be executed concurrently across multiple devices, thereby effectively distributing both memory overhead and computational workload.
@@ III-A2 Tensor Parallelism
Megatron-LM TP divides tensors into one-dimensional segments. While this approach reduces model parameters stored on each device to $\frac{1}{N_{d}}$, it still needs to store complete intermediate activations on each device. Additionally, forward and backward computations of each transformer layer involve 4 all-reduce, causing significant communication overhead. To address this, researchers have proposed multidimensional tensor partitioning methods [86], [87], [88], which further reduce memory overhead on each device and mitigate the frequent communication issues of 1D TP [89]. In practice, TP is typically applied within nodes rather than across nodes to prevent communication from becoming a bottleneck for model training efficiency [90].
III-A3 Pipeline Parallelism
Pipeline parallelism, also known as inter-layer parallelism, involves partitioning the model layers horizontally into multiple stages. Each stage consists of consecutive layers, which are assigned to different devices for sequential execution. Unlike TP, which divides intra-layer tensor calculations, PP partitions the model into stages, where each device transfers activations and gradients between adjacent stages during forward and backward passes. PP not only reduces memory demands and computational workloads on each device but also has less communication overhead. However, a major challenge associated with PP is device idle time (often referred to as pipeline bubbles). While asynchronous PP schemes [91], [92] can eliminate pipeline bubbles, they have no guarantee of model convergence [93]. As synchronous PP schemes continue to improve, their performance and efficiency are comparable with asynchronous PP, making them a more favorable option. In the following paragraphs, we will discuss in detail several common PP schemes.
In vanilla PP, data can be passed to the next device only after the preceding device completes its calculation, due to the sequential dependence in data flow. This leads to pipeline bubbles, resulting in a waste of resources and diminishing overall efficiency of the system.
GPipe [94] splits each input batch into several micro-batches, which are input to the model sequentially. After completing the processing of each micro-batch, each device immediately passes its results to the next device and begins to process the next micro-batch. This approach significantly reduces the pipeline bubbles, and finally, each device uses the accumulated gradients to perform synchronous parameter updates. To further enhance memory efficiency, GPipe incorporates gradient checkpointing [79], which retains only a subset of intermediate activations and recalculates the remaining activations during backpropagation. However, GPipe still faces the issue of high memory consumption for intermediate activations.
The GPipe scheduling involves completing the forward passes for all micro-batches before executing the backward passes (F-then-B). Let $N_{s}$ denote the number of stages in the model, $N_{d}$ denote the number of devices, and $N_{m}$ denote the number of micro-batches. In the GPipe schedule, each device must retain intermediate activations of $N_{m}$ micro-batches. To reduce memory consumption, researchers introduced a one forward pass followed by one backward pass pattern (1F1B) [95], [96]. This scheduling strategy effectively reduces the number of on-the-fly micro-batches by executing the backward pass in advance. With 1F1B schedule, the pipeline bubble is equal to that of GPipe, and the activations of micro-batches required to store on each device decrease from $N_{m}$ to $N_{d}$ (where $N_{m}$ is typically greater than $N_{d}$).
The aforementioned PP schemes assume that the number of pipeline stages $N_{s}$ and the number of devices $N_{d}$ are equal, with each device possessing one stage. Narayanan et al. [95] split the stages based on the 1F1B schedule, allowing each device to have multiple stages. For instance, in Fig. 4, an 8-layer model is partitioned into 4 stages, where each stage consists of two consecutive layers, and each device handles one stage (device 1 possesses L1 and L2, device 2 possesses L3 and L4, etc.). If the model is divided into 8 stages, each device will be responsible for the computation of 2 stages. This interleaved 1F1B schedule performs a finer-grained scheduling of forward and backward computations, reducing pipeline bubbles and enhancing efficiency. However, the trade-off for this reduction in pipeline bubbles includes increased communication and higher activation memory usage.
Figure 4: Device placements of 1F1B and interleaved 1F1B. Interleaved 1F1B partitions model into finer-grained stages and assign each device with multiple stages to reduce pipeline bubbles.
In training iteration, the computational cost of backpropagation is greater than that of forward propagation (generally estimated to be about twice as much), resulting in longer computation time for backpropagation. This is the primary contributor to the existence of pipeline bubbles. Qi et al. [97] suggest splitting the backward pass into the calculation of the gradients of activations (denoted as $B$), and the calculation of the gradients of parameters (denoted as $W$). They found that the $B$ pass is independent of the $W$ pass within the current layer and only depends on the $B$ pass from the previous layer. Theoretically, one could complete $B$ passes of all layers before performing $W$ passes for each layer. This allows for the design of more efficient pipeline schedules. One of the proposed pipeline schedules (ZB-H1) is similar to 1F1B, but it offers the flexibility to schedule $W$ after the corresponding $B$ within the same stage, effectively filling bubbles in the schedule. Additionally, in terms of activation memory usage, ZB-H1 ensures that peak memory usage on all devices does not exceed 1F1B. If the memory limit is adjusted to accommodate peak memory usage exceeding 1F1B, along with the optimizer post-validation strategy proposed by the author, zero bubble scheduling (ZB-H2) can be achieved. ZB-H2 fills additional $F$ during the warm-up phase to fill bubbles prior to the initial $B$, followed by a rearrangement of $W$ passes at the tail to completely eliminate all bubbles in pipeline schedule.
PP typically has the issue of high activation memory usage, and the 1F1B pipeline schedules mentioned above have the problem of memory imbalance. Regarding this, Qi et al. [93] proposed an analytic framework to decompose existing pipeline schedules into building blocks and found that peak memory usage of a pipeline schedule is highly related to the lifespan of the building blocks. Based on this, a series of V-Shape building blocks are proposed in the article, which have more balanced peak memory on each device compared to interleaved 1F1B.
In the distributed training of large-scale deep learning models, PP is a commonly used technique. Compared to TP, it only communicates between adjacent stages and has lower communication overhead. Therefore, it is usually applied between nodes to scale up model size [95].
III-A4 Sequence Parallelism
Supporting long sequence lengths in large-scale models is crucial for applying AI to scientific endeavors. Many scientific challenges inherently involve complex analysis of large-scale, high-dimensional data, frequently presented as long sequences. Long-sequence issues are prevalent in fields including structural biology, chemistry, drug development, and atmospheric science. For instance, many proteins consist of hundreds or even thousands of amino acid residues, where the sequence context of each residue can be essential to its 3D structure. Long-sequence support enhances the ability to capture non-local interactions among residues, thereby improving the quality of predictions. Additionally, complex molecules or chemical reactions often necessitate long sequence representations, including molecular formulas, SMILES, reaction pathways, etc. Climate prediction models require analyzing long-term historical and global spatial data, represented in time series or grid formats. Despite the importance of long-sequence support, its implementation presents challenges. Processing long sequences demands significant increases in computational power and memory, with the memory usage of attention mechanisms rising quadratically with sequence length.
Jacobs et al. proposed Ulysses [99], a method that divides the input along sequence dimension onto all devices and conducts all-to-all communication on the Query (Q), Key (K), and Value (V) matrices before performing attention computation. This ensures that each device operates with disjoint attention heads, allowing for distributed attention calculation. Once the calculations are complete, another all-to-all is performed to get the results and restore the state that inputs were divided along sequence dimension. Ulysses SP and Megatron-LM TP share similarities in the distributed calculation of self-attention, both concerning limitations on the number of attention heads. A key advantage of Ulysses is that the communication volume for a single device decreases linearly as SP degree increases, while in Megatron-LM, it is independent of TP degree.
Liu et al. [100] developed a ring-style distributed attention computation method that optimizes resource usage by conserving a portion of the Q, K, and V blocks. During calculations, each device exchanges K and V blocks with its neighboring devices, enabling it to perform computations using blockwise self-attention and feedforward networks (FFN) [101]. In contrast to the approach taken by Li et al. [98], this method can eliminate communication overhead by picking an proper chunk size and overlapping communication with computation.
III-B Mixed Precision Training

### v2

@@ 2.2.2 Medicine
LLMs have demonstrated significant potential in the medical and clinical fields. The applications of such medical models are numerous, including medical knowledge retrieval, assistance in clinical diagnoses, consultations, and personalized medical services.
@@
Singhal et al. [3] found that the benchmark used to evaluate the clinical knowledge of the model is limited. To solve it, they proposed MultiMedQA, a benchmark composed of seven medical Q&A datasets. Further, they fine-tuned the command of PaLM [28] and got Flan-PaLM [26], which shows excellent performance in multiple-choice questions. Moreover, they suggested instruction prompt tuning, and the obtained Med-PaLM model can further adapt to medical field.
@@
Tu et al. [5] aim to further explore the potential of generalist models (those capable of performing multiple tasks without fine-tuning, such as GPT-3, PaLM, and GPT-4) in medical field. Utilizing the language and multimodal foundation models PaLM 540B and PaLM-E 562B [51], the researchers develop Med-PaLM M. They also designed a multimodal biomedical benchmark called MultiMedBench. As a biomedical generalist model, Med-PaLM M demonstrates strong performance across all tasks on MultiMedBench, matching or exceeding the capabilities of task-specific models.
@@
Currently, most LLMs utilized in medical field are either closed source or have relatively small model sizes. Chen et al. [39] believe that while models designed for generalist tasks typically perform well across tasks, their performance in certain tasks falls short compared to models tailored for those specific purposes. They developed Meditron-70B through continued pre-training of Llama-2 and fine-tuned it for particular tasks, resulting in superior performance compared to all open-source generalist and medical LLMs across all medical benchmarks.
@@
HuatuoGPT [54] aims to assist doctors with diagnosis and treatment in Chinese scenarios, it combines the dialogue capabilities of ChatGPT with the expertise of real-world physicians. By employing instruction fine-tuning and reinforcement learning techniques, it exploits distilled instruction data generated by ChatGPT alongside actual response data from doctors. The upgraded version, HuatuoGPT-II [56], successfully passed the Chinese National Pharmacist Licensure Examination in October 2023 and nearly all medical qualification exams, highlighting its robust capabilities within the Chinese medical landscape.
2.2.3 Biomedicine
@@
2.2.4 Chemistry
The application of LLM in the field of chemistry covers multiple aspects, including molecular design, molecular property prediction, and chemical reaction prediction. Additionally, LLMs can be used to analyze and summarize chemical literature, providing researchers with efficient knowledge retrieval and decision-making support.
@@ 2.2.4 Chemistry
Irwin et al. pre-trained BART [63] on a dataset containing a large number of simplified molecular line entry system (SMILES) strings to capture the structural relationships and potential properties between chemical molecules, resulting in the Chemformer model [9]. Chemformer is capable of both sequence-to-sequence tasks (e.g. reaction prediction, molecular optimization), and discriminative tasks (e.g. molecular property prediction, biological activity analysis). It achieves or exceeds the performance of state-of-the-art (SOTA) models on multiple benchmarks while significantly reducing convergence time.
@@
The objective of molecular design and generation tasks is to create new molecules with specific functions or properties. Traditional methodologies depend heavily on the expertise of specialists and costly computational simulations. Fang et al. introduced a molecular generative pre-trained language model called MolGen [7], which has enhanced its comprehension of chemical structures and semantics through a two-stage pre-training strategy. This strategy encompasses molecular syntax and semantic learning, utilizing self-sacrificing embedded strings along with domain-agnostic prefix tuning. Frey et al. [6] conducted an in-depth study on the neural scaling behavior of deep chemical models in relation to dataset and model size, aiming to explore how model performance is enhanced with increased resource investment. They designed and trained a generative pre-trained Transformer model named ChemGPT with over 1B parameters for small molecule generation. By varying both the model and dataset sizes by several orders of magnitude, they demonstrated an improvement in pre-training loss performance as the scale increased.
2.2.5 Meteorology
@@
Over the past decade, the advancement of high-performance computing devices has led to significant progress in numerical weather prediction (NWP), enhancing the accuracy of forecasts for daily weather, extreme weather events, and climate change. However, as the growth of computing power slows and the complexity of physical models increases, the limitations of traditional methods have become more pronounced. Recently, the rapid progress in DL has emerged as a promising avenue for addressing these challenges.
@@
The Pangu-Weather model [10, 11], proposed by Bi et al., employs the 3D Earth-Specific Transformer (3DEST) architecture to process 3D climate data, which is a variant of Swin Transformer [65, 66]. It is the first AI-driven approach to surpass the accuracy of NWP methods. Notably, Pangu-Weather achieves an inference speed of 1.4 seconds on a single GPU, making it 10000× faster than the previous most accurate NWP method (operational IFS [74]). While the accuracy of 3D models can surpass 2D models, this enhancement comes with a considerable increase in memory requirements, leading to limited network depth.
@@
The inherent uncertainty in weather forecasting is inevitable, and the degree of this uncertainty, along with cumulative errors, increase as the forecast period lengthens. In response to this challenge, Chen et al. introduced FuXi [12], a cascaded model architecture based on U-Transformer, which effectively reduce cumulative errors and improves both forecast accuracy and duration. The model pre-training was conducted on a cluster of 8 Nvidia A100 GPUs.
@@
Inspired by the concept of foundational models [75], Nguyen et al. designed and trained a basic model named ClimaX [13], aimed at effectively adapting to general tasks related to the Earth’s atmosphere. ClimaX can be trained on heterogeneous datasets to solve diverse climate and weather tasks. Remarkably, even with pre-training conducted at lower resolutions and with limited computational resources, ClimaX surpasses previous data-driven models in weather and climate prediction benchmarks. The pre-training used 80 NVIDIA V100 32GB GPUs. Subsequently, Bodnar et al. introduced Aurora [14], a atmospheric foundation model designed to excel in various prediction tasks, particularly in regions with limited data or extreme weather conditions. After training on diverse weather and climate datasets, Aurora can effectively adapt to new atmospheric prediction tasks by learning a universal representation of atmospheric dynamics. The model was trained on 32 GPUs with 1 batch size per GPU.
2.2.6 Geoscience
As LLMs demonstrate success in multiple domains, expanding their capabilities to Earth science becomes a natural progression. The geoscience field, with its extensive and complex datasets, demands efficient and domain-aware LLM solutions.
K2 [68] pioneers the application of LLMs in geoscience by extending LLaMA-7B through domain-specific pre-training and expert instruction tuning using the GeoSignal dataset. Evaluations on the GeoBench benchmark reveal K2’s superior performance over general LLMs in both factual recall and reasoning tasks within the geoscience domain.
Building on this trend, GeoGalactica [69] further scales geoscience LLMs by specializing in a 30B-parameter model through continued pre-training of Galactica [70] on 65B tokens, followed by instruction fine-tuning. The training relies on Megatron-LM framework and utilizes a computing cluster consisting of 512 nodes, with each node equipped with 4 hygon DCU accelerator. GeoGalactica achieves SOTA performance on various geoscience benchmarks and expert-validated tasks, underscoring the importance of domain-adaptive scaling and memory-efficient training for advancing AI-driven Earth system research.
Complementing these efforts, JiuZhou [71] introduces a geoscience foundation model developed through a two-stage pre-adaptation pre-training method on the large-scale JiuZhou-Corpus. Despite utilizing a limited two-GPU setup, JiuZhou effectively scales training with long-context inputs and large batch sizes, achieving competitive performance on the GeoBench benchmark and outperforming GPT-3.5 in various geoscientific tasks. This work further illustrates the growing importance of memory-efficient training strategies for advancing LLMs in Earth science.
2.2.7 Memory Challenges in Scaling Transformers for Scientific Applications
@@
Despite the progress achieved in applying transformer models across various domains, the continuous increase in memory overhead during model training remains a pressing concern that requires attention. We identify the underlying causes of this issue from the following key factors:
@@
  • The demand for handling diverse data structures. In scientific fields, the use of high-resolution images and multimodal data also significantly contributes to increased activation memory consumption during large model training.
{forest}
@@
3.2.1), system [Data Parallelism, system [Pytorch DDP [94], ZeRO [95], FSDP [47], system-end, text width=12cm] ] [Tensor Parallelism, system [Megatron-LM [96], 2D-TP [97], 2.5D-TP [98], 3D-TP [99], system-end, text width=12cm]] [Pipeline Parallelism, system [GPipe [100], 1F1B [101, 102, 103], Interleaved 1F1B [102], Zero Bubble [104, 105], system-end, text width=12cm]] [Sequence Parallelism, system [Ring Self-Attention [106], Megatron SP [93], Ulysses [107], Ring Attention [108], system-end, text width=12cm] ] ] [Offloading (§3.2.2), system [SwapAdvisor [109], ZeRO-Offload [110], ZeRO-Infinity [111], system-end] ] ] [Hardware-Software Co-Optimization (§3.3), hardware [FlashAttention [112], FlashAttention-2 [113], FlashAttention-3 [114], MPress [115], hardware-end, text width=14.35cm] ] ]
Figure 2: Overview of memory-efficient training techniques for transformers.
3 Memory-Efficient Training Techniques of Transformers
LLMs are evolving rapidly and find extensive applications across various domains. However, the training of these models imposes significant demands on computational resources and memory. In this context, exploring memory-efficient training strategies is critical to enabling broader accessibility and supporting larger model sizes. While LLMs have gained considerable popularity in scientific fields and have led to noteworthy results, our analysis suggests that many existing works employ only a limited subset of available memory optimization techniques. As shown in Table 1, most works adopt one or more common memory-saving techniques (e.g., mixed precision training, DP), but more advanced strategies remain underutilized. This may limit the ability to train larger or more efficient models. In the rest of this section, we will delve into various memory-efficient training techniques specific to transformer-based models, as illustrated in Fig. 2.
3.1 Algorithm
A variety of algorithmic techniques have been proposed to address memory challenges. These techniques optimize memory usage while maintaining training efficiency and model accuracy. This section categorizes and analyzes algorithmic strategies for memory optimization in transformer models. Table 2 provides an overview of various algorithm-level memory-efficient training methods, summarizing their characteristics and typical use cases.
3.1.1 Compression
Compression techniques reduce memory usage by minimizing the precision or redundancy of data representations during training, while aiming to retain comparable performance. These methods are widely used in resource-constrained scenarios.
  1. i.
    Mixed Precision Training